# A Concise Lyapunov Analysis of Nesterov’s Accelerated Gradient Method

### v2

@@ 2 Preliminaries
We recall some basic definitions and results for unconstrained smooth optimization, which can be found in [8, Chapter 2]. We use $\left|x\right|$ to denote the Euclidean norm of $x\in\mathbb{R}^{n}$ and $x\cdot y$ to denote the inner product of $x,y\in\mathbb{R}^{n}$. The gradient of a differentiable function $f:\,\mathbb{R}^{n}\rightarrow\mathbb{R}$ is denoted by $\nabla f$.
@@
$\left|f(x)-f(y)\right|\leq L\left|x-y\right|,$
@@
The following well-known inequality for $L$-smooth functions will be used later.
@@
[8, Lemma 1.2.3] If $f:\,\mathbb{R}^{n}\rightarrow\mathbb{R}$ is $L$-smooth, then
@@ 4 Strongly convex functions
The proof is inspired by the estimating sequence argument in [8], which treats both general (weakly) and strongly convex functions in a unified fashion. While estimating sequences are powerful, their construction is inductive, and the analysis of the algorithm becomes entangled with its design. Lyapunov analysis is well known for effectively analyzing the stability of dynamical systems [4]. Since optimization algorithms can be viewed as discrete-time dynamical systems, it is not surprising that Lyapunov functions can be used to analyze their convergence [10, 14]. In particular, [14] provides a comprehensive study of Lyapunov analysis for accelerated methods in optimization as discretizations of continuous-time dynamical systems. Despite this comprehensive work, we believe a concise and direct Lyapunov analysis can be beneficial in revealing the convergence behavior of algorithm (5). We encourage interested readers to contrast the analysis in [14], particularly Proposition 13, as well as the proof using estimating sequences in [8, Section 2.2] with the proof of Theorem 8 presented here.
@@ 5 Conclusions
This work was partially supported by the NSERC of Canada and the CRC program. The author is grateful to Nicolas Boumal for identifying a typo in an earlier version that caused an issue in the proof of Theorem 5.

### v3

@@ 1 Introduction
The use of Lyapunov functions (sometimes called potential functions) for the analysis of optimization algorithms has been popularized over the years by various authors. In addition to [18] mentioned above, readers are referred to [2, Section 5], [6, Section 4], [4, Section 4], and references therein. Readers are also referred to dissipative theory, which is closely related to Lyapunov analysis, in the context of Nesterov’s accelerated method [7]. We will add several remarks throughout the paper to point readers to closely related work.
@@ 2 Preliminaries
We recall some basic definitions and results for unconstrained smooth optimization, which can be found in [12, Chapter 2]. We use $\left|x\right|$ to denote the Euclidean norm of $x\in\mathbb{R}^{n}$ and $x\cdot y$ to denote the inner product of $x,y\in\mathbb{R}^{n}$. The gradient of a differentiable function $f:\,\mathbb{R}^{n}\rightarrow\mathbb{R}$ is denoted by $\nabla f$.
@@
$\left|\nabla f(x)-\nabla f(y)\right|\leq L\left|x-y\right|,$
@@
The following well-known inequality for $L$-smooth functions (also known as the descent lemma) will be used later.
@@
[12, Lemma 1.2.3] If $f:\,\mathbb{R}^{n}\rightarrow\mathbb{R}$ is $L$-smooth, then
@@ 4 Strongly convex functions
The proof is inspired by the estimating sequence argument in [12], which treats both general (weakly) and strongly convex functions in a unified fashion. While estimating sequences are powerful, their construction is inductive, and the analysis of the algorithm becomes entangled with its design. Lyapunov analysis is well known for effectively analyzing the stability of dynamical systems [8]. Since optimization algorithms can be viewed as discrete-time dynamical systems, it is not surprising that Lyapunov functions can be used to analyze their convergence [14, 18]. In particular, [18] provides a comprehensive study of Lyapunov analysis for accelerated methods in optimization as discretizations of continuous-time dynamical systems. Despite this comprehensive work, we believe a concise and direct Lyapunov analysis can be beneficial in revealing the convergence behavior of algorithm (5). We encourage interested readers to contrast the analysis in [18], particularly Proposition 13, as well as the proof using estimating sequences in [12, Section 2.2] with the proof of Theorem 8 presented here.
Remark 10.
Dissipativity theory for Nesterov’s accelerated methods has been investigated in [7]. A key benefit of this framework is that it enables the construction of Lyapunov functions, either numerically or analytically, by solving small semidefinite programs. In [7], the authors provided analytical solutions to LMIs for both the general and strongly convex cases of Nesterov’s method; see [7, Theorem 9] and [7, Theorem 4], respectively. While their results are more general and parametrized by LMI solutions, Theorems 5 and 8 in this paper provide concise and self-contained proofs for the basic forms of Nesterov’s scheme. These may be of interest to readers seeking simplified derivations or alternative insights.
@@ 5 Conclusions
This work was partially supported by the NSERC of Canada and the CRC program. The author is grateful to Nicolas Boumal for identifying a typo in an earlier version that caused an issue in the proof of Theorem 5. The author is also grateful to reviewers of an earlier version of this paper, who provided additional references.